OpenStudy (anonymous):

Solve it! (5-x)1/2 =x+1 First answer gets a Medal!

mathslover (mathslover):

is it : \(\sf (5-x) \times \cfrac{1}{2} = x + 1 \) or something else?

OpenStudy (anonymous):

Free medal for you maths :)

OpenStudy (anonymous):

yes @mathslover

OpenStudy (anonymous):

doteki, what have you done?

mathslover (mathslover):

Okay , multiply 2 both sides. What do you get?

OpenStudy (anonymous):

I my suppose to multiply 5-x to 1/2?

mathslover (mathslover):

Yes but in order to get away from 1/2 , multiply 2 both sides : \( (5-x) \times \cfrac{1}{2} \color{blue}{\times 2} = (x+1) \color{blue}{\times \cfrac{1}{2}}\)

OpenStudy (anonymous):

alright let me try to solve that part, give me a moment

mathslover (mathslover):

Yeah, sure, take your time.

OpenStudy (anonymous):

yeah i really can't solve it I don't understand how to solve this equation. I'm stuck at 2.5-1/2x

OpenStudy (anonymous):